### v1

@@ Abstract
Context:The upcoming Square Kilometer Array (SKA) will set a new standard regarding data volume generated by an astronomical instrument, which is likely to challenge widely adopted data analysis tools that scale inadequately with the data size.
Aims:This study aims to develop a new source detection and characterization method for massive radio astronomical datasets by adapting modern deep-learning object detection techniques. These approaches have proved their efficiency on complex computer vision tasks, and we seek to identify their specific strengths and weaknesses when applied to astronomical data.
Methods:We introduce YOLO-CIANNA, a highly customized deep-learning object detector designed specifically for astronomical datasets. This paper presents the method and describes all the low-level adaptations required to address the specific challenges of radio-astronomical images. We demonstrate this method’s capabilities using simulated 2D continuum images from the SKA Observatory (SKAO) Science Data Challenge 1 (SDC1) dataset.
Results:Our method outperforms every other published result on the specific SDC1 dataset. Using the SDC1 metric, we improve the challenge-winning score by +139% and the score of the only other post-challenge participation by +61%. Our catalog has a detection purity of 94% while detecting 40 to 60 % more sources than previous top-score results with a total of almost 680000 properly detected sources. The trained model can also be forced to reach 99% purity in post-process and still detect 10 to 30% more sources than the other top-score methods. Our method is efficient at low signal-to-noise ratio and exhibits strong characterization accuracy. It is also capable of real-time detection, with a peak prediction speed of 500 images of $512{\times}512$ pixels per second on a single GPU.
Conclusions:YOLO-CIANNA achieves state-of-the-art detection and characterization results on the simulated SDC1 dataset. This is encouraging regarding its potential capability over observational data from SKA precursors. The method is open source and included in the wider CIANNA framework. We provide scripts to train and apply this method to the SDC1 dataset in the CIANNA repository.
YOLO-CIANNA: Galaxy detection with deep learning in radio data
I. A new YOLO-inspired source detection method applied to the SKAO SDC1
D. Cornu 11    P. Salomé 11    B. Semelin 11    A. Marchal 2233    J. Freundlich 44    S. Aicardi 55    X. Lu 66    G. Sainton 11    F. Mertens 11    F. Combes 1177    C. Tasse 8899
(Received …, …; accepted …, …)
@@ 1 Introduction
Modern astronomical instruments generate ever-increasing data volumes, following the need for better resolution, sensitivity, and larger wavelength coverage. Astronomical datasets are often highly dimensional and require precise encoding of the measurements due to a high dynamic range. In addition, it is often necessary to preserve the raw data due to iterative improvement of the analysis pipelines. Radio-astronomy is strongly affected by the explosion of data volumes, especially regarding giant radio interferometers. In particular, the upcoming Square Kilometer Array (SKA, Braun et al., 2019) is expected to have an unprecedented real-time data rate and to produce a remarkable amount of stored science data products with around 700 PB of archived data per year. This instrument is foreseen to have the necessary sensitivity to set constraints on the cosmic dawn and the epoch of reionization and to trace the evolution of astronomical objects over cosmological times. With such volume and complexity of data, some classical analysis methods and tools employed in radio astronomy for decades start to exhibit scaling limits.
@@
In this context, the SKA Observatory (SKAO) started the organization of recurrent Science Data Challenges (SDCs) to gather astronomers from the international community around simulated datasets that resemble future SKA data products. The objective is to evaluate the suitability of existing analysis methods and encourage the development of new ones. It is also an opportunity for astronomers to get familiar with the nature of such datasets and to gain experience in their exploration.
@@
The first edition, SDC1 (Bonaldi et al., 2021), focused on a source detection and characterization task over simulated continuum radio images at different frequencies and integration times. We show a cutout from one of the SDC1 images in Fig. 1, illustrating the source density and the high dynamical range. Source-finding is a common task in astronomy and is often the first analysis to be done on a newly acquired image product. It is already performed by a variety of classical methods, for example, SExtractor (Bertin & Arnouts, 1996), SFIND (Hopkins et al., 2002), CUTEX (Molinari et al., 2011), BLOBCAT (Hales et al., 2012), AEGEAN (Hancock et al., 2018), DUCHAMP (Whiting, 2012), PyBDSF (Mohan & Rafferty, 2015), PROFOUND (Robotham et al., 2018). The obtained source catalogs can then be augmented with characterization information and used as primary data for subsequent analyses. This task is strongly affected by the increase in volume and dimensionality, making it a good proxy to evaluate the upcoming data handling challenges.
@@
In the past decade, we observed an explosion in the machine learning (ML) methods usage in all fields, including astronomy and astrophysics (Huertas-Company & Lanusse, 2023). One of the advantages of ML methods is their good performance and efficiency scaling with data size and dimensionality. There is a considerable variety of ML approaches, so we only focus here on methods based on the deep artificial neural networks formalism (LeCun et al., 2015). Deep learning approaches have been extensively used for computer vision tasks, including the leading object detection application (Russakovsky et al., 2015; Everingham et al., 2010; Lin et al., 2014). While detection models have been extensively used in other domains for several years, they are not yet widely adopted in the astronomical community.
@@
Deep learning object detection methods are usually separated into three families (Zhao et al., 2018). The first one represents segmentation models. Their main advantage is identifying which pixels belong to a given object type (semantic segmentation). Their main drawback is their symmetric structure (encoder and decoder) and the amount of work to be done near the image resolution, making them compute-intensive. They can also be used as a convenient structure for denoising tasks. This family is mainly represented by the U-Net (Ronneberger et al., 2015) method. Due to their proximity with classical source detection approaches, they have been employed for a variety of astronomical applications (e.g., Akeret et al., 2017; Vafaei Sadr et al., 2019; Lukic et al., 2019; Paillassa et al., 2020; Bianco et al., 2021; Makinen et al., 2021; Sortino et al., 2023; Håkansson et al., 2023).
@@
The second family corresponds to the region-based detectors. They are often based on multi-stage neural networks that split the detection task into a region proposal step and a detection refinement step. They are the most employed for mission-critical tasks due to their accuracy. While faster than segmentation methods, the high detection accuracy models are compute-intensive due to the multi-stage process. This family is mainly represented by the R-CNN method (Girshick et al., 2013) and all its derivatives (e.g., Fast R-CNN, Faster R-CNN). Examples of astronomical applications with these methods are more limited, but it is increasing (e.g., Wu et al., 2019; Jia et al., 2020; Lao et al., 2021; Yu et al., 2022; Sortino et al., 2023). There is a special variation of these methods that combines the region-based detection formalism with a mask prediction used to perform instance segmentation. They are mainly represented by the Mask R-CNN method (He et al., 2017), which is also increasingly used in astronomy (e.g., Burke et al., 2019; Farias et al., 2020; Riggi et al., 2023; Sortino et al., 2023). We note that region-based methods are commonly combined with feature pyramid network (Lin et al., 2016), which helps represent multiple scales in the detection task.
@@
The last family consists of regression-based detectors, which are mostly based on single-stage neural networks. These methods are compute-efficient and often used for real-time object detection. They are mainly represented by the YOLO method and its sub-versions Redmon et al. (2015); Redmon & Farhadi (2016, 2018), but we can also cite SSD (Liu et al., 2015). There have been a few astronomical applications, mostly in the visible domain (González et al., 2018; He et al., 2021; Wang et al., 2021; Grishin et al., 2023; Xing et al., 2023).
@@
We highlight that methods based on transformers (Vaswani et al., 2017) are now common in computer vision (Carion et al., 2020), and astronomical applications are just starting to be published (Gupta et al., 2024; He et al., 2023). We also note that some methods include deep learning parts in more classical source detection tools, which can improve the detection purity or the source characterization (e.g., Tolley et al., 2022). More references regarding deep learning methods for source detection can be found in Sortino et al. (2023) an Ndung’u et al. (2023).
@@
This is the first paper of a series that aims to present a new source detection and characterization method called YOLO-CIANNA that was developed and used in the context of MINERVA’s (MachINe lEarning for Radioastronomy at Observatoire de Paris) team participation in the SDC2 (Hartley et al., 2023), enabling it to achieve first place. This first paper describes the method and presents its application over simulated 2D continuum images from the SDC1 dataset. A second paper will present an application over simulated 3D cubes of HI emission using the SDC2 dataset. The series will then continue applying the method to observational data from several SKA precursors.
The primary objective of this first paper is to describe the YOLO-CIANNA method, which is done in Sect. 2, and to present how we adapted it to account for the specific challenges of astronomical source detection. In Sect. 3, we present the SDC1 dataset, composed of comprehensive 2D images, and expose how we used it to construct a benchmark to evaluate our method’s detection and characterization capabilities. In Sect. 4, we present the detection result of our method and do a detailed analysis of the source catalog we obtained over the SDC1. We use these results to highlight the strengths and weaknesses of our detector, which are then discussed in Sect. 5. We also added three significant Appendix sections. The first one, Appendix A, presents the differences between our YOLO-CIANNA method and the classical YOLO implementation. In Appendix B, we present how the classical network architecture associated with YOLO would perform on the SDC1. And finally, Appendix A, presents an alternative training area definition for the SDC1.
@@
Figure 1: Cutout of 512 square pixels in the SDC1 560 MHz 1000h simulated field. Minimum and maximum cutting values are those used for our object detector, but the image dynamic is not altered.
@@ 2 Method
Our method was inspired by the You Only Look Once (YOLO, Redmon et al., 2015; Redmon & Farhadi, 2016, 2018) approach, a regression-based deep learning object detector. While region-based approaches like R-CNN (Girshick et al., 2013) are often considered the most accurate object detectors, regression-based methods present a straightforward single network architecture, making them more compute-efficient at a given detection accuracy. Both families can reach state-of-the-art accuracy depending on implementation details and architecture design. YOLO-like methods are usually preferred for real-time detection applications. In this context, our choice of exploring a YOLO-inspired regression-based approach was driven by i) fewer implementation constraints, ii) a strong emphasis on compute performance considering the upcoming data volume of radio-astronomical surveys, and iii) the single network regression-based structure on which it is easier to add more predictive capabilities.
In this section, we present the main design and properties of our custom object detection method along with necessary general concepts about object detection for non-expert readers. Despite being depicted for an astronomical application, our method remains suitable for general-purpose object detection (Appendix A.8). For clarity, we describe the whole method from scratch, which includes aspects from the classical YOLO implementation and our dedicated modifications. The added or modified elements in comparison to the three first classical versions from the original author, YOLO-V1 (Redmon et al., 2015), YOLO-V2 (Redmon & Farhadi, 2016), and YOLO-V3 (Redmon & Farhadi, 2018) will be mentioned. Still, the more technical and in-depth justifications for these changes are presented in Appendix A. Even though the modifications we brought to the YOLO algorithm are substantial, we still refer to our approach as YOLO-CIANNA in this paper for the sake of simplicity.
The implementation was made inside the custom high-performance deep learning framework CIANNA [^1] CIANNA is open source and freely accessible through GitHub https://github.com/Deyht/CIANNA. The version used in this paper corresponds to the 1.0 release. DOI:xx.xxxx/xxxxx.xxxxxxx (Convolutional Interactive Artificial Neural Networks by/for Astrophysicists). The implementation and usage details can be found on the CIANNA wiki pages. For reproducibility purposes, we provide example scripts for training and applying the method to the SDC1 dataset in the CIANNA git repository.
To ease the understanding of the technical parts of the paper for readers unfamiliar with ML terminology, we list a few technical terms we use and the associated descriptions we have for them. The most common ML terms are not defined but can be found in any proper ML textbook or review (LeCun et al., 2015).
@@ 2.1 Bounding boxes for object detection
Our method uses a fully convolutional neural network (CNN) structure to construct a mapping from a 2D input image to a regular output grid of detection units. Each output grid cell represents a small area of the input image with a size that depends on the ratio between the input and the output grid resolutions. Each grid cell is tasked to detect all possible objects whose center is located inside the input region it represents. To characterize an object, we rely on the bounding box formalism that encodes an object as a four-dimension vector composed of the box center and its size ($x,y,w,h$), which are the quantities that the detection units must predict. Our method belongs to the supervised learning approaches, so it relies on a training set composed of images with a list of all the visible objects to be detected. Each object can be encoded as a target bounding box that the detector will be tasked to predict using only information from the input image. This can be done through an optimization process, also called learning, which is an iterative process that aims at minimizing a loss function $\mathcal{L}$, also called an error function, that compares the target boxes with the predicted boxes at the current step. This loss should encompass all the object properties to be predicted. To ease the method description, we first write an abstract loss as
$\mathcal{L}=\mathcal{L}_{pos}+\mathcal{L}_{size}+\mathcal{L}_{prob}+\mathcal{L}_{obj}+\mathcal{L}_{class}+\mathcal{L}_{param}.$ (1)
The aim of the Sects. 2.1 to 2.4 is to describe all of the loss sub-parts. Our complete detailed loss function is presented in Sect. 2.7 with Eq. 14.
For now, we only describe the case of a single box prediction per grid cell. The more realistic case of multiple objects per grid cell is presented in Sect. 2.5. To represent a bounding box, each grid cell must predict a 4-element vector ($o^{x},o^{y},o^{w},o^{h}$) that maps to the box’s geometric properties following
@@
Each grid cell is only tasked to position the object center inside its dedicated area, which is obtained using two sigmoid-activated values ($o^{x},o^{y}$). The position of the grid cell in the full image must then be added to obtain the real position of the object, which is expressed by the ($g_{x},g_{y}$) coordinates. The object size is obtained by an exponential transform of the predicted values ($o^{w},o^{h}$) that acts as a scaling on a pre-defined size prior ($p_{w},p_{h}$). This is equivalent to an anchor-box formalism (Ren et al., 2015) as discussed in Sect. 2.5. The corresponding bounding box construction on the output grid is illustrated in Fig. 2.
@@
Figure 2: Illustration of the YOLO bounding box representation. The different quantities correspond to Eqs. 2 to 5. The dashed black box corresponds to the theoretical prior ($o_{w}=o_{h}=0$), while the red box represents the scaled network predicted size.
@@
With this formalism, it is possible to construct a network output layer with a $g_{w},g_{h},4\right>$ grid that is capable of positioning and scaling a bounding box for each grid cell. For each prediction-target pair, we use a sum-of-square error to compute the output loss function for center coordinates and sizes (Sect. 2.7). The error is not computed on the sigmoid-activated positions but on the raw output for the sizes after target conversion using $\hat{o}^{w}=\log(w/p_{w})$ and $\hat{o}^{h}=\log(h/p_{h})$. This results in the following loss terms
$\mathcal{L}_{pos}$ $=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}\left((o^{x}_{i}-\hat{o}^{x}_{i})^{2}+(o^{y}_{i}-\hat{o}^{y}_{i})^{2}\right),$ (6)
$\mathcal{L}_{size}$ $=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{match}\left((o^{w}_{i}-\hat{o}^{h}_{i})^{2}+(o^{h}_{i}-\hat{o}^{h}_{i})^{2}\right),$ (7)
@@
where the hat values represent the target for the corresponding predicted value, the sum over $i$ represents all the grid cells with $N_{g}=g_{w}{\times}g_{h}$, and $\mathds{1}_{i}^{match}$ is a mask to identify the predicted boxes that have an associated target box (Sect. 2.6). The grid cells that do not contain any object have no contribution to these loss terms. All these elements follow the classical YOLO formalism. We discuss the possible limitations of using bounding boxes to describe astronomical objects in Sect. 5.2.2.
@@
We emphasize that nothing prevents the size of the predicted box from being larger than the area mapped by a grid cell up to the size of the full image. Each grid cell receives information from a large area corresponding to the backbone network receptive field. The receptive fields of nearby grid cells usually overlap, but a target box center can only lie in one grid cell. Due to the fully convolutional structure required for our method, each grid cell represents a localized prediction using identical weights. It is equivalent to having a single detector that scans different regions of the same image but in a more efficient way from a network architecture standpoint. This approach is equivalent to what is done starting with YOLO-V2 but differs from the one introduced in YOLO-V1. More details about the effect of the fully convolutional architecture and the corresponding output grid encoding are provided in Appendix A.1 and A.2.

### v2

@@ Abstract
Context. The upcoming Square Kilometer Array (SKA) will set a new standard regarding data volume generated by an astronomical instrument, which is likely to challenge widely adopted data-analysis tools that scale inadequately with the data size.
Aims. The aim of this study is to develop a new source detection and characterization method for massive radio astronomical datasets based on modern deep-learning object detection techniques. For this, we seek to identify the specific strengths and weaknesses of this type of approach when applied to astronomical data.
Methods. We introduce YOLO-CIANNA, a highly customized deep-learning object detector designed specifically for astronomical datasets. In this paper, we present the method and describe all the elements introduced to address the specific challenges of radio astronomical images. We then demonstrate the capabilities of this method by applying it to simulated 2D continuum images from the SKA observatory Science Data Challenge 1 (SDC1) dataset.
Results. Using the SDC1 metric, we improve the challenge-winning score by +139% and the score of the only other post-challenge participation by +61%. Our catalog has a detection purity of 94% while detecting 40 to 60 % more sources than previous top-score results, and exhibits strong characterization accuracy. The trained model can also be forced to reach 99% purity in post-process and still detect 10 to 30% more sources than the other top-score methods. It is also computationally efficient, with a peak prediction speed of 500 images of $512{\times}512$ pixels per second on a single GPU.
Conclusions. YOLO-CIANNA achieves state-of-the-art detection and characterization results on the simulated SDC1 dataset and is expected to transfer well to observational data from SKA precursors.
@@ 1 Introduction
Modern astronomical instruments generate ever-increasing data volumes in response to the need for better resolution, sensitivity, and wider wavelength intervals. Astronomical datasets are often highly dimensional and require precise encoding of the measurements due to a high dynamic range. Also, it is often necessary to preserve the raw data so they can be re-analyzed with new versions of the processing pipelines. Radio-astronomy is strongly affected by this explosion of data volumes, especially regarding giant radio interferometers. The upcoming Square Kilometer Array (SKA, Braun et al., 2015) is expected to have an unprecedented real-time data-production rate, and will provide 700 PB of archived data per year (Scaife, 2020). This instrument is foreseen to have the necessary sensitivity to set constraints on the cosmic dawn and the epoch of reionization and to trace the evolution of astronomical objects over cosmological times. Faced with data of such volume and complexity, some classical analysis methods and tools employed in radio astronomy will struggle to keep up due to their limited scalability.
@@
In this context, the SKA Observatory (SKAO) began organizing recurrent Science Data Challenges (SDCs) in order to gather astronomers from the international community around simulated datasets that resemble future SKA data products. The objective is to evaluate the suitability of existing analysis methods and encourage the development of new ones. It is also an opportunity for astronomers to become familiar with the nature of such datasets and to gain experience in their exploration.
@@
The first edition, SDC1 (Bonaldi et al., 2021), focused on a source detection and characterization task in simulated continuum radio images at different frequencies and integration times. Figure 1 shows a cutout from one of the SDC1 images, illustrating the source density and high dynamic range. Source-finding is a common task in astronomy and is often the first analysis carried out on a newly acquired image product; it is already performed by a variety of classical methods, such as Source-Extractor (Bertin & Arnouts, 1996), SFIND (Hopkins et al., 2002), CUTEX (Molinari et al., 2011), BLOBCAT (Hales et al., 2012), DUCHAMP (Whiting, 2012), SELAVY (Whiting & Humphreys, 2012), AEGEAN (Hancock et al., 2018), PyBDSF (Mohan & Rafferty, 2015), PROFOUND (Robotham et al., 2018), PySE (Carbone et al., 2018), CAESAR (Riggi et al., 2019), and CERES (Lucas et al., 2019). The obtained source catalogs can then be augmented with characterization information and used as primary data for subsequent studies. This task is especially affected by increases in volume and dimensionality, making it a good probe of the upcoming data-handling challenges.
@@
The past decade has seen a rapid increase in the use of machine learning (ML) methods in all fields, including astronomy and astrophysics (Huertas-Company & Lanusse, 2023). One of the advantages of ML methods is their superior scaling with data size and dimensionality. There are a considerable variety of ML approaches, but we focus here on methods based on deep artificial neural networks (LeCun et al., 2015). These approaches have been extensively used for computer vision tasks, including object detection in everyday-life images (Russakovsky et al., 2015; Everingham et al., 2010; Lin et al., 2014). While detection models have been used in other domains for several years, they are not yet widely adopted in the astronomical community.
@@
Deep-learning object-detection methods are usually separated into three families (Zhao et al., 2019): segmentation models, region-based detectors, and regression-based detectors. The main advantage of the segmentation models is their ability to identify all the pixels that belong to a given class or even individual objects. They can also be used as a convenient structure for denoising tasks. Their main drawback is their symmetric structure (encoder and decoder) and the high level of expressivity required at near-input resolution, making them computationally intensive for high-resolution images. This family is mainly represented by the U-Net (Ronneberger et al., 2015) method. Given their proximity with classical source detection approaches, they have been employed for a variety of astronomical applications (e.g., Akeret et al., 2017; Vafaei Sadr et al., 2019; Lukic et al., 2019; Paillassa et al., 2020; Bianco et al., 2021; Makinen et al., 2021; Sortino et al., 2023; Håkansson et al., 2023).
@@
The second family, the region-based detectors, mostly comprise multi-stage neural networks that split the detection task into a region-proposal step and a detection-refinement step. They are the most popular method for mission-critical tasks due to their accuracy. While faster than segmentation methods, these high-detection-accuracy models are computationally intensive due to the multi-stage process. This family is mainly represented by the R-CNN method (Girshick et al., 2013) and all its derivatives (e.g., Fast R-CNN, Faster R-CNN). Examples of astronomical applications with these methods are more limited (e.g., Wu et al., 2019; Jia et al., 2020; Lao et al., 2021; Yu et al., 2022; Sortino et al., 2023). There is also a subfamily that combines the region-based detection formalism with a mask prediction used to perform instance segmentation. This subfamily is mainly represented by the Mask R-CNN method (He et al., 2017), which is increasingly used in astronomy (e.g., Burke et al., 2019; Farias et al., 2020; Riggi et al., 2023; Sortino et al., 2023). We note that region-based methods are commonly combined with some flavor of pyramidal feature hierarchy construction (Lin et al., 2017), which helps represent multiple scales in the detection task.
@@
The last family, the regression-based detectors, are often based on single-stage neural networks, making them computationally efficient. Therefore, they are frequently used for real-time object detection. The most popular regression-based detector method is the You Only Look Once (YOLO) method and its descendants (Redmon et al., 2016; Redmon & Farhadi, 2017, 2018), but we can also cite the Single Shot Detector (SSD) method (Liu et al., 2015). There have only been a few astronomical applications of regression-based detectors, mainly in the visible domain (González et al., 2018; He et al., 2021; Wang et al., 2021; Grishin et al., 2023; Xing et al., 2023).
@@
We highlight that methods based on transformers (Vaswani et al., 2017) are now common in computer vision (Carion et al., 2020), and astronomical applications are just starting to be published (Gupta et al., 2024; He et al., 2023). We also note that some methods include deep learning parts in more classical source detection tools, which can improve the detection purity and the source characterization (e.g., Tolley et al., 2022). Further references regarding deep learning methods for source detection can be found in Sortino et al. (2023) and Ndung’u et al. (2023).
@@
The present paper is the first of a series, the aim of which is to present a new source detection and characterization method called YOLO-CIANNA. It was developed and used in the context of the MINERVA (MachINe lEarning for Radioastronomy at Observatoire de Paris) team participation in the SDC2 (Hartley et al., 2023), where we obtained first place. The primary objective of this first paper is to provide a complete description of the method and to justify several design choices regarding the specific properties of astronomical images. We then illustrate its capability by presenting the results of its application to simulated 2D continuum images from the SDC1 dataset. A second paper will present an application to simulated 3D cubes of HI emission using the SDC2 dataset. The series will then continue by applying the method to observational data from surveys of several SKA precursors.
The present paper is organized as follows:
  • In Sect. 2, we present the YOLO-CIANNA method in a complete and comprehensive manner, targeting readers unfamiliar with deep-learning object detectors. Details regarding the most complex aspects of the method and the in-depth differences from a classical YOLO implementation are given in Appendix A along with details adapted to readers familiar with these families of methods.
  • In Sect. 3, we present the SDC1 dataset, which is composed of comprehensive 2D images, and describe how we used it as a benchmark to evaluate the detection and characterization capabilities of YOLO-CIANNA.
  • In Sect. 4, we present the detection results of our method, as well as a detailed analysis of the detection catalog we obtained from the SDC1.
  • In Sect. 5, we use these results to highlight the strengths and weaknesses of our source-detector and also discuss the impact of some design choices of the SDC1. We then elaborate on how this new approach could be applied to real observational data from SKA precursor instruments.
  • There are four Appendices. In Appendix A, we detail the differences between our method and the classical YOLO implementation, as well as benchmarks over classical computer vision datasets. In Appendix B, we evaluate the performance of a classical YOLO backbone architecture on the SDC1 and compare it to our custom backbone. In Appendix C, we present an alternative training area definition for the SDC1. In Appendix D, we evaluate whether performing the detection and characterization using a single unified network is beneficial or detrimental to detection-only performance.
@@
Figure 1: Cutout of 512 square pixels in the SDC1 560 MHz 1000h simulated field. Minimum and maximum cutting values are those used for our object detector, and the displayed intensity is the raw image flux.
@@ 2 Method
Our method finds its inspiration in the classical YOLO (Redmon et al., 2016; Redmon & Farhadi, 2017, 2018) object detector, but it is also very similar to the SSD method (Liu et al., 2015), both regression-based deep learning object detector. While region-based approaches like R-CNN (Girshick et al., 2013) are often considered the most accurate object detectors, regression-based methods present a straightforward single network architecture, making them more computationally efficient at a given detection accuracy. Both families can reach state-of-the-art accuracy depending on implementation details and architecture design, but regression methods are usually preferred for real-time detection applications. In this context, our choice to design a regression-based approach was driven by i) fewer implementation constraints, ii) a strong emphasis on computational performance considering the data volume of current and future radio-astronomical surveys, and iii) the single-stage regression-based network structure on top of which it is easier to add more predictive capabilities.
In this section, we present the main design of our method. Despite being depicted for an astronomical application, our method remains suitable for general-purpose object detection (Appendix A.8). Instead of listing all the subtle differences with various descendant versions of the classical YOLO approach or other detectors, we describe the main components of our method from scratch in a comprehensive manner. The description of the most technical parts of the method is done in the dedicated Appendix A, which also discusses the differences with the classical YOLO approach. As a result, the main method description should remain accessible to readers unfamiliar with deep-learning object detectors. Even though our method differs significantly from the YOLO algorithm in some critical aspects, we chose to refer to our approach as YOLO-CIANNA for simplicity and as a legacy for its inspiration.
The implementation was made inside the custom high-performance deep learning framework CIANNA (Cornu, 2024b) [^1] CIANNA is open source and freely accessible through GitHub https://github.com/Deyht/CIANNA. The version used in this paper corresponds to the V-1.0 release 10.5281/zenodo.12806325. The implementation and usage details can be found on the CIANNA wiki pages. For reproducibility purposes, we provide example scripts for training and applying the method to the SDC1 dataset in the CIANNA git repository.
To ease the understanding of the technical parts of the paper, we list a few ML-specific terms and the descriptions we have for them. The most common ML terms are not redefined as they can be found in any ML textbook or review (LeCun et al., 2015).
@@ 2.1 Bounding boxes for object detection
Our method uses a fully convolutional neural network (CNN) to construct a mapping from a 2D input image to a regular output grid. Each output grid cell represents a fixed area of the input image with a size that depends on the reduction factor of the chosen CNN backbone. Each grid cell is tasked to detect all possible objects whose center is located inside the input region it represents. To characterize an object, we rely on the bounding box formalism that encodes an object as a four-dimension vector composed of the box center and its size ($x,y,w,h$). The grid cells are tasked to predict these quantities. As a supervised learning approach, our method relies on a training set composed of images associated with a list of visible objects to be detected. Each object can be encoded as a target bounding box that the CNN is tasked to predict from the raw input image. A training phase is used to optimize the network parameters to minimize a loss function $\mathcal{L}$ that compares the target boxes with the predicted boxes from all grid cells at the current step. This loss should encompass all the object properties to be predicted. To ease the method description, we first write an abstract loss as
$\mathcal{L}=\mathcal{L}_{\textrm{pos}}+\mathcal{L}_{\textrm{size}}+\mathcal{L}_{\textrm{prob}}+\mathcal{L}_{\textrm{obj}}+\mathcal{L}_{\textrm{class}}+\mathcal{L}_{\textrm{param}}.$ (1)
The aim of the Sects. 2.1 to 2.4 is to describe all of the loss subparts. Our complete detailed loss function is presented in Sect. 2.7 with Eq. 14.
For now, we only describe the case of a single box prediction per grid cell. The more realistic case of multiple objects per grid cell is presented in Sect. 2.5. We consider that the grid is composed of $g_{w}$ columns and $g_{h}$ lines and that each grid cell is represented by its coordinate in the grid ($g_{x}$,$g_{y}$). To represent a bounding box, a grid cell predicts a 4-element vector ($o^{x},o^{y},o^{w},o^{h}$) that maps to the geometric properties of the box following
@@
Each grid cell is only tasked to position the object center inside its dedicated area using two sigmoid-activated values ($o^{x},o^{y}$). The position of the grid cell in the image ($g_{x},g_{y}$) is added to obtain the relative global position of the object. These coordinates must then be multiplied by the mapped width and height of a grid cell, corresponding to the reduction factor of the backbone network, to obtain pixel coordinates in the input image. Object size is obtained by an exponential transform of the predicted values ($o^{w},o^{h}$) that acts as a scaling on a predefined size prior ($p_{w},p_{h}$), which can be expressed in pixels directly. This is equivalent to an anchor-box formalism (Ren et al., 2015) as discussed in Sect. 2.5. The corresponding bounding box construction on the output grid is illustrated in Fig. 2.
@@
Figure 2: Illustration of the YOLO bounding box representation following Eqs. 2 to 5. The dashed black box represents the theoretical prior ($o_{w}=o_{h}=0$), while the red box is the scaled predicted size.
@@
With this formalism, it is possible to construct a detector with an output of size $g_{w},g_{h},4\right>$ that can position and scale one bounding box per output grid cell. For each prediction-target pair, we use a sum-of-square error to compute the loss function for center coordinates and sizes (Sect. 2.7). The error is not computed on the sigmoid-activated positions but on the raw output for the sizes after target conversion using $\hat{o}^{w}=\log(w/p_{w})$ and $\hat{o}^{h}=\log(h/p_{h})$. This results in the following loss terms
$\mathcal{L}_{\textrm{pos}}$ $=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left((o^{x}_{i}-\hat{o}^{x}_{i})^{2}+(o^{y}_{i}-\hat{o}^{y}_{i})^{2}\right),$ (6)
$\mathcal{L}_{\textrm{size}}$ $=\sum_{i=0}^{N_{g}}\mathds{1}_{i}^{\textrm{match}}\left((o^{w}_{i}-\hat{o}^{w}_{i})^{2}+(o^{h}_{i}-\hat{o}^{h}_{i})^{2}\right),$ (7)
@@
where the hat values represent the target for the corresponding predicted value, the sum over $i$ represents all the grid cells with $N_{g}=g_{w}{\times}g_{h}$, and $\mathds{1}_{i}^{\textrm{match}}$ is a mask to identify the predicted boxes that have an associated target box (Sect. 2.6). The grid cells that do not contain any object have no contribution to these loss terms. We discuss the possible limitations of using bounding boxes to describe astronomical objects in Sect. 5.2.2.
@@
A common misconception about grided detection is that the predicted size can only be as large as a grid element, which is wrong. The predicted size can be as large as necessary up to the size of the full image. Each grid cell receives information from a large area corresponding to the backbone network receptive field. The receptive fields of nearby grid cells usually overlap, but a target box center can only lie in one grid cell, hence the attribution of the detection to a single grid cell. Due to the fully convolutional structure required for our method, each grid cell performs a localized prediction using identical weights but using a different subpart of the image as input. More details about the effect of the fully convolutional architecture and the output grid encoding are provided in Appendix A.1 and A.2.